OpenStudy (anonymous):

HELP!! A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days. f(n) = 12(1.03)^n Part A: When the scientist concluded his study, the height of the plant was approximately 16.13 cm. What is a reasonable domain to plot the growth function?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@texaschic101

OpenStudy (solomonzelman):

I got disconnected, I am really sorry _:(

OpenStudy (anonymous):

its fine

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf f(n)=12(1.03)^{n} }\) is the function, where f(n) represents the height of the plant, and n represents the day(s). you know that the height of the plant is 16.13, plug that instead of f(n) \(\LARGE\color{blue}{ \bf 16.13=12(1.03)^{n} }\) can you solve for n? ( Just curious, is this algebra 1 ? )

OpenStudy (anonymous):

yes, it is

OpenStudy (solomonzelman):

then I don't think you will be able to get an EXACT answer, however you can estimate it, but hardly. If I show you the exact solution I am not sure if you get it, but do you want to try it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I will try it

OpenStudy (anonymous):

10.006299982393

OpenStudy (anonymous):

?

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf 16.13 =12(10.3)^{n}}\) divide both sides by 12, and I get, \(\LARGE\color{blue}{ \bf 1.34 =(10.3)^{n}}\) now, I am hitting both sides with a log, \(\LARGE\color{red}{ \bf log~1.34 =log~(10.3)^{n}}\) now I am simplifying the log, \(\LARGE\color{red}{ \bf log~1.34 =n ~log~(10.3)}\) \(\LARGE\color{red}{ \bf log~1.34~~/~~log~(10.3) =n ~}\) \(\LARGE\color{red}{ \bf log~_{(10.3) }1.34=n ~}\) \(\LARGE\color{red}{ \bf n=0.13 ~}\)

OpenStudy (solomonzelman):

Well, I think the domain would then be approaching zero.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

What does the y-intercept of the graph of the function f(n) represent?

OpenStudy (solomonzelman):

the y-intercept, is when n equals zero. So, if n=0 that means that no days have passed, or in other words, it is the initial height of the plant.

OpenStudy (anonymous):

uh... isn't it 1.03 and not 10.3?